# Axion Astrophysics

### v1

@@ Abstract
journal: Physics Reportsjournal: Physics Reports
@@ 1 Introduction
Figure 1: Stellar axion spectra at Earth (number flux in the upper panel, and energy flux in the lower panel) vs axion energy $\omega_{a}$, when considering axions coupled to photons and produced by different sources: the Sun (total flux in dashed line) via Primakoff [37, 38] – black line, and in case of a magnetic field longitudinal plasmon (LP) conversion [39] and transverse plasmon (TP) conversion [40] – grey bands; a close-by red supergiant like Betelgeuse [26] (blue band) at a distance $d=197$ pc, and a $18~{}M_{\odot}$ Galactic SN at a distance of 10 kpc [41] (green band). There are represented also the diffuse fluxes from extra-galactic main-sequence stars [42] [Diffuse Main-Sequence Axion Background (DMSAB) – magenta band] and from extra-galactic SNe, the Diffuse Supernova Axion Background (DSAB) [43, 28] (orange band). It is assumed an axion-photon coupling $g_{a\gamma}=10^{-11}~{}{\rm GeV}^{-1}$ and an axion mass $m_{a}\ll{\rm eV}$.
@@
So, why another review? The objective of this review is not to replace but to update and, in some cases, complete the previous monoghraps. The field of axion astrophysics has experienced a substantial development in the recent years. Practically all axion bounds from stellar evolution have been updated. Furthermore, the methodology itself has changed in recent years. Several new astrophysical mechanisms and observables, previously largely ignored, are now considered in the attempt to extract axion properties in the most reliable way. Finally, the increasingly precise astrophysical observations ask for more accurate numerical simulations and for a careful analysis of the uncertainties that affect observations and simulations. Thus, we consider it timely to provide a comprehensive and updated review on the state-of-the-art axion astrophysics, which includes the modern developments and might provide a good reference for both theorists and experimentalists interested in axion physics. We hope this will complete the list of recently published excellent reviews on axion cosmology [47], axion experiments [48, 49], and axion models [50]. Finally, since astrophysical arguments have for decades guided experimental axion searches – a fact that is getting more and more relevant nowadays, since new axion experiments are finally reaching the potential to probe parameter regions of astrophysical interest [51, 22] – we hope that our review could provide also a good reference also for the upcoming experimental searches.
@@
Other interesting aspects of axion astrophysics, including very recent results from X-ray observations [52, 53, 54], the quite studied topic of very-high-energy astrophysics [55, 56, 57, 58, 59, 60, 61], the conversions of dark matter axions in the magnetic field of astrophysical objects like NS [62, 63, 64, 65] and the Sun [66, 67], or axion constraints from the modified equations of state in compact stars [68, 69] or black-hole superradiance [70, 71] have not been included in our review.
@@ 2 Stellar evolution and feebly-interacting particles
In this Section we will revise the physics principle ruling the stellar evolution and how to account for the role of FIPs. In Sec. 2.1 we present the differential equations describing the stellar evolution. In Sec. 2.2 we describe which are the main sources and sinks of energy in stellar interiors. In Sec. 2.3 we characterize the radiative energy transport, while in Sec. 2.4 we consider the effect of convection. In Sec. 2.5 we describe the impact of FIPs on the equations of stellar evolution. Finally, in Sec. 2.6 we present a brief overview of the stellar evolution, introducing different types of stars we will refer to in the next Sections. We remark that, for pedagogical reasons, in this section we employ CGS units while in the rest of the report we adopt natural units. A discussion on the different unit systems is provided in Appendix A.
@@ 2.1 Stellar structures: a minimal set of equations
They provide an appropriate description of the internal structure of the majority of the stars, when the effects of rotation and magnetic fields can be neglected. The first equation represents the equilibrium between gravity and pressure. Here, $m_{r}$ is the mass contained within the sphere of radius $r$ and $P$ is the pressure at $r$. The second equation describes the mass conservation. Here, $\rho$ is the density at $r$. Hence, the third is the energy conservation equation. $L_{r}$ is the luminosity at $r$, i.e., the radiated electromagnetic power trough the spherical surface with radius $r$, while the three $\varepsilon$ coefficients in the right hand side of this equation represent the rates of energy generation or sink (per unit mass) as due to the release of gravitational energy ($\varepsilon_{\rm grav}$) and nuclear energy ($\varepsilon_{\rm nucl}$), and the energy loss due to the production of thermal neutrinos ($\varepsilon_{\nu}$). Then, the fourth equation describes the energy transport that is explicitly connected to the temperature gradient. The nabla function ($\nabla=\frac{d\log T}{d\log P}$) depends on the specific transport mechanism, radiation, convection or conduction. Finally, Eq. (6) describes the variations of the chemical composition due to the occurrence of nuclear reactions or bulk motion of material, such as that induced by thermal convection. Here, $X_{i}$ are the mass fractions and $n$ is the number of isotopes. When coupled to the equation of state (EoS), $P=f(\rho,T,\mu)$, and appropriate expressions for the energy generation and energy sink rates, $\varepsilon_{\rm grav,nucl,\nu}$, numerical solutions of Eqs. (2)–(6) provide values of $4+n$ dependent variables (e.g., $r$, $L$, $P$, $T$ and the $n$$X_{i}$) as a function of the mass coordinate, i.e., the Lagrangian variable $m_{r}$, and the time $t$. The quantity $\mu$ is the mean molecular weight. In case of a fully ionized plasma, it is
@@
From the equations described so far, it is evident that a star evolves to compensate the energy losses, and it implies a change in the internal structure. In addition, also mass loss processes should be considered. In practice, the stellar structure is continuously adjusted in order to maintain thermal and hydrostatic equilibrium. In principle, the $n+4$ differential equations [Eqs. (2))–(6))] are strictly coupled and should be simultaneously solved. This is a bit time consuming, so that, in most cases, the solution of the first four equations and that of the $n$ equations [Eq. (6)] are obtained in two separate steps. Firstly, the $n$ equations describing the evolution of the chemical composition are solved by keeping the temperature and the density profiles fixed during a time step, and then, the 4 equations describing the physical structure are solved by taking the chemical composition fixed. This is a good approximation when the physical quantities, like $T$ or $L$, change slowly compared to the variations of the chemical composition. In other words, the evolution timescale of the physical structure must be much longer than the nuclear burning timescale. When dynamical instability, such as convection, are at work, also the mixing timescale should be small compared to the timescales of hydrostatic relaxation. In general, this condition is quite well fulfilled by core-H and core-He burning stars, but they are easily violated during more advanced phases of stellar evolution. In all the cases the majority of the stellar evolution codes use an implicit method to solve the stellar structure equation, as firstly proposed by Ref. [73]. So, differences in the models found in the extant literature are mainly due to differences in the input physics [EoS, radiative opacity, nuclear network and nuclear reaction rates, energy loss rates]. The numerical accuracy (mesh point spacing and time steps) may also produces discrepancies. Modern stellar evolution codes use adaptive mesh and time step algorithms to ensure accurate solutions all along a complete evolutionary sequence.
@@ 2.2 Sources and sinks of energy
where $N_{A}$ is the Avogadro’s number and $\langle\sigma v\rangle$ is the rate of the reaction. In the case of a reaction between charge particles, it is
@@ 2.5.3 Constraining the physics nature of FIPs
The previous argument has been applied to several classes of stars (Sun, RG, He-burning stars, WD, NS, etc…) and several FIP candidates. A general introduction is provided in Ref. [15]. In this work we will focus on the production and the impact of axions on stellar evolution. In the next sections, a brief overview of stellar evolution for different stellar masses is provided, focusing, in particular, on the evolutionary phases in which a larger impact on a FIP production is expected.
@@ 2.6.1 The virial theorem
Making use of Eq. (3), let us rewrite the hysdrostatic equilibrium equation [Eq. (2)]
@@ 2.6.4 H and He burning
In low-mass stars ($M<2~{}{\rm M}_{\odot}$) electron degeneracy develops in the He-rich core. This occurrence has two major effects. Firstly the electron conductivity increases, contributing to an efficient energy redistribution within the core. As a result, an almost isothermal temperature profile develops within the core. On top of that, a plasma neutrino emission is more effective in the innermost layers, where the density is higher. As a consequence of this neutrino cooling, the maximum temperature moves out of the center (see Fig. 4). Then, an off-center He ignition occurs, which causes a thermonuclear runaway known has He-flash. In practice, the release of nuclear energy by the triple-$\alpha$ reactions causes an increase of the temperature not counterbalanced by an increase of the pressure. This is a consequence of the degenerate equation of state, for which the electron pressure mainly depends on their density, while it weakly depends on $T$. As shown in Fig. 4, up to $10^{10}$ $L_{\odot}$ of nuclear power are attained by the first and more powerful flash, which is followed by a series of weaker thermonuclear runaways, until the electron degeneracy is removed in the whole core. This occurrence marks the beginning of the quiescent He burning and it is called the Zero Age Horizontal Branch (ZAHB).
@@ 2.6.6 The final destiny of low-mass and intermediate-mass stars
During this phase, the star climbs the so-called AGB. The compact C-O core is surrounded by a thin He-rich mantle and a more external and very expanded H-rich envelope. The stellar radius becomes very large, i.e., a few hundreds R. Two burning shells are alternatively actives. For most of the time the shell-H burning provides enough energy to replace the energy lost by radiation. Meanwhile the He-rich mantel is accreted with the ashes of the H-burning, until it reaches a critical mass. Then, recursive thermal instability powered by the periodic activation of the He-burning takes place (thermal pulses). The discovery of these thermal instabilities [100] had a great relevance in our understanding of the chemical evolution of the Universe. About half of the elements heavier than iron are produced by low-mass AGB stars through a slow-neutron-capture nucleosynthesis process or s process [101]. Moreover, AGB stars are among the major sources of dust in a galaxy like the Milky Way. The AGB mass loss rate may be as large as $10^{5}$ M/yr, and the wind from these stars may be as fast as $10^{6}$ cm/s. Owing to this huge mass loss, the H-rich envelope is rapidly eroded until a bare core remains. These stars end their life as C-O WD.
Figure 10: Evolution of the central temperature and density for stars having masses between 1 and 9 ${\rm M}_{\odot}$ and nearly solar composition [96]. The three principal evolutionary phases are labeled, namely: core-H burning (H), core-He burning (He) and the AGB. Note, in particular, the expansion occurring after the He ignition, as due to the release of nuclear energy, and the cooling of the core during the final part of the evolutionary tracks, when stars climb the AGB. This occurrence is a consequence of the production of plasma-neutrinos, which extracts energy from the stellar core. The red-dashed curves represent constant degeneracy parameter lines, namely: $\phi=\frac{\mu}{kT}=-10,0,10,100$, left to right. Here $\mu$ is the electron chemical potential that, in the $T=0$ limit, coincides with the Fermi energy. This curves have been calculated assuming an electron fraction $Y_{e}=0.5$, which is appropriate for the He-rich core of RGB stars as well as for the C-O core of AGB stars. Low-mass stars ($M<2.5~{}{\rm M}_{\odot}$) develop a partial degeneracy of the electron ($\phi>0$) during the first ascent of the RGB. Later on, during the AGB phase, the electron degeneracy develops again, in all stars with mass lower than $\sim 9~{}{\rm M}_{\odot}$. Eventually, the yellow area delimits the region where electrons are degenerate and relativistic. Note that the more massive AGB stars ($M>4~{}{\rm M}_{\odot}$) enter in this region when the plasma-neutrino cooling is more efficient.
@@ 3.1 Axions and axion-like particles
Several extensions of the SM of particle physics predict a hidden sector of new particles interacting very weakly with SM particles. Often, such scenarios include very light particles, for example as Nambu-Goldstone bosons of global symmetries [10, 9, 102, 103, 104, 105], or as modes associated with compactified extra-dimensions [106, 107, 108]. The study of these additional light modes goes under the name of the low-energy frontier of particle physics [18] and has become one of the most studied branches of modern particle physics. Indeed, a general interest in the FIPs has emerged in recent years, particularly since they offer explanations to several important cosmological and astrophysical problems (see Refs. [17, 16] for a recent reviews).
@@ 00 KSVZ
Although different models can provide very different phenomenological results, in general, axions are very constrained by the requirement to solve the strong CP problem (see Ref. [124] for more details). Here we just summarize a few aspects  [^7] The interested reader may find more information in Ref. [50].:
@@
Figure 12: Contours of the axion energy-loss rates per unit mass, $\varepsilon_{a}$, in erg g-1s-1, for a pure He plasma for ordinary stars, while only nucleons and electron fraction $Y_{e}=0.3$ for supernovae (SNe) and neutron stars (NSs). Different lines represent different channels, as shown in the legend. The Primakoff process (blue lines) is calculated for $g_{a\gamma}=10^{-11}$eV-1. The Bremsstrahlung (black lines) and Compton (green lines) processes are calculated for $g_{ae}=10^{-13}$. The onset of the degeneracy region is visible in the bending of the bremsstrahlung contours. The shaded regions are indicative of the central temperature and density of the Sun [142], Red Giant Branch (RGB) stars, horizontal branch (HB) stars and SNe are also shown. In the case of HB and RGB, these are the results of a numerical simulation of a 0.8 $M_{\odot}$ model as obtained with the FuNS code [88]. With the same code also the Betelgeuse model is obtained [26]. Regarding degenerate stars, we show the properties of WD L19-2 and two NSs, representing the case of a young NS, J172054 with an age of $\sim 650$ yr, and an old one, J1856 with an age of $4\times 10^{5}$ yr [36]. The thickness of the lines has no significance, except for SNe, where roughly represents the plausible range of density and temperature.

### v2

@@ Abstract
journal: Physics Reports
@@ 1 Introduction
Figure 1: Stellar axion spectra at Earth (number flux in the upper panel, and energy flux in the lower panel) vs axion energy $\omega_{a}$, when considering axions coupled to photons and produced by different sources: the Sun (total flux in dashed line) via Primakoff [37, 38] – black line, and in case of a magnetic field longitudinal plasmon (LP) conversion [39] and transverse plasmon (TP) conversion [40] – gray bands; a close-by red supergiant like Betelgeuse [26] (blue band) at a distance $d=197$ pc, and a Galactic SN at a distance of 10 kpc with a progenitor of $18~{}M_{\odot}$ [41] (green band). There are represented also the diffuse fluxes from extra-galactic main-sequence stars [42] [Diffuse Main-Sequence Axion Background (DMSAB) – magenta band] and from extra-galactic SNe, the Diffuse Supernova Axion Background (DSAB) [43, 28] (orange band). It is assumed an axion-photon coupling $g_{a\gamma}=10^{-11}~{}{\rm GeV}^{-1}$ and an axion mass $m_{a}\ll{\rm eV}$.
@@
So, why another review? The objective of this review is not to replace but to update and, in some cases, complete the previous monoghraps. The field of axion astrophysics has experienced a substantial development in the recent years. Practically all axion bounds from stellar evolution have been updated. Furthermore, the methodology itself has changed in recent years. Several new astrophysical mechanisms and observables, previously largely ignored, are now considered in the attempt to extract axion properties in the most reliable way. Finally, the increasingly precise astrophysical observations ask for more accurate numerical simulations and for a careful analysis of the uncertainties that affect observations and simulations. Thus, we consider it timely to provide a comprehensive and updated review on the state-of-the-art axion astrophysics, which includes the modern developments and might provide a good reference for both theorists and experimentalists interested in axion physics. We hope this will complete the list of recently published excellent reviews on axion cosmology [47], axion experiments [48, 49], and axion models [50]. Finally, since astrophysical arguments have for decades guided experimental axion searches – a fact that is getting more and more relevant nowadays, since new axion experiments are finally reaching the potential to probe parameter regions of astrophysical interest [51, 22] – we hope that our review could provide also a good reference for the upcoming experimental searches.
@@
Other interesting aspects of axion astrophysics, including very recent results from X-ray observations [52, 53, 54], the quite studied topic of very-high-energy astrophysics [55, 56, 57, 58, 59, 60, 61], the conversions of dark matter axions in the magnetic field of astrophysical objects like NS [62, 63, 64, 65], (primordial) black holes [66] and the Sun [67, 68], or axion constraints from the modified equations of state in compact stars [69, 70] or black-hole superradiance [71, 72] have not been included in our review.
@@ 2 Stellar evolution and feebly-interacting particles
In this Section we will revise the physics principle ruling the stellar evolution and how to account for the role of FIPs. In Sec. 2.1 we present the differential equations describing the stellar evolution. In Sec. 2.2 we describe the main sources and sinks of energy in stellar interiors. In Sec. 2.3 we characterize the radiative energy transport, while in Sec. 2.4 we consider the effect of convection. In Sec. 2.5 we describe the impact of FIPs on the equations of stellar evolution. Finally, in Sec. 2.6 we present a brief overview of the stellar evolution, introducing different types of stars we will refer to in the next Sections. We remark that, for pedagogical reasons, in this section we employ CGS units while in the rest of the report we adopt natural units. A discussion on the different unit systems is provided in Appendix A.
@@ 2.1 Stellar structures: a minimal set of equations
They provide an appropriate description of the internal structure of the majority of the stars, when the effects of rotation and magnetic fields can be neglected. The first equation represents the equilibrium between gravity and pressure. Here, $m_{r}$ is the mass contained within the sphere of radius $r$ and $P$ is the pressure at $r$. The second equation describes the mass conservation. Here, $\rho$ is the density at $r$. The third is the energy conservation equation. $L_{r}$ is the luminosity at $r$, i.e., the radiated electromagnetic power through the spherical surface with radius $r$, while the three $\varepsilon$ coefficients in the right hand side of this equation represent the rates of energy generation or sink (per unit mass) as due to the release of gravitational energy ($\varepsilon_{\rm grav}$) and nuclear energy ($\varepsilon_{\rm nucl}$), and the energy loss due to the production of thermal neutrinos ($\varepsilon_{\nu}$). Then, the fourth equation describes the energy transport that is explicitly connected to the temperature gradient. The nabla function ($\nabla=\frac{d\log T}{d\log P}$) depends on the specific transport mechanism, radiation, convection or conduction. Finally, Eq. (6) describes the variations of the chemical composition due to the occurrence of nuclear reactions or bulk motion of material, such as that induced by thermal convection. Here, $X_{i}$ are the mass fractions and $n$ is the number of isotopes. When coupled to the equation of state (EoS), $P=f(\rho,T,\mu)$, and appropriate expressions for the energy generation and energy sink rates, $\varepsilon_{\rm grav,nucl,\nu}$, numerical solutions of Eqs. (2)–(6) provide values of $4+n$ dependent variables (e.g., $r$, $L$, $P$, $T$ and the $n$$X_{i}$) as a function of the mass coordinate, i.e., the Lagrangian variable $m_{r}$, and the time $t$. The quantity $\mu$ is the mean molecular weight. In case of a fully ionized plasma, it is
@@
From the equations described so far, it appears that a star evolves because of energy losses and changes in the internal chemical composition due, for example, to nuclear reactions. In particular, time derivatives appear explicitly in Eqs. (6), which describe variations of the chemical composition, and implicitly in Eq. (4), where the left-hand side represents the outward energy fluence, while the coefficients $\epsilon$ in the right-hand side represent rates of energy production or consumption. Also, mass-loss or mass-accretion processes should be taken into account. In practice, the stellar structure is continuously adjusted in order to maintain thermal and hydrostatic equilibrium. In principle, the $n+4$ differential equations [Eqs. (2))–(6))] are strictly coupled and should be solved simultaneously. This is a somewhat time consuming, and in most cases the solution of the first four equations and that of the $n$ equations [Eq. (6)] are obtained in two separate steps. Firstly, the $n$ equations describing the evolution of the chemical composition are solved by keeping the temperature and the density profiles fixed during a time step, and then, the 4 equations describing the physical structure are solved by taking the chemical composition fixed. This is a good approximation when physical quantities, such as $T$ or $L$, change slowly compared to variations in the chemical composition. In other words, the evolution timescale of the physical structure must be much longer than the nuclear burning timescale. When dynamical instability, such as convection, are at work, also the mixing timescale should be short compared to the timescales of hydrostatic relaxation. In general, this condition is quite well fulfilled by core-H and core-He burning stars, but they are easily violated during more advanced phases of stellar evolution. In all the cases the majority of the stellar evolution codes use an implicit method to solve the stellar structure equation, as firstly proposed by Ref. [74]. So, differences in the models found in the extant literature are mainly due to differences in the input physics [EoS, radiative opacity, nuclear network and nuclear reaction rates, energy loss rates]. The numerical accuracy (mesh point spacing and time steps) may also produces discrepancies. Modern stellar evolution codes use adaptive mesh and time step algorithms to ensure accurate solutions all along a complete evolutionary sequence.
@@ 2.2 Sources and sinks of energy
where $N_{A}$ is the Avogadro’s number and $\langle\sigma v\rangle$ is the rate of the reaction. In the case of a reaction between charged particles, it is
@@ 2.5.3 Constraining the physics nature of FIPs
The previous argument has been applied to several classes of stars (Sun, RG, He-burning stars, WD, NS, etc…) and several FIP candidates. A general introduction is provided in Ref. [15]. In this work we will focus on the production and the impact of axions on stellar evolution. In the next sections, a brief overview of stellar evolution for different stellar masses is provided, focusing, in particular, on the evolutionary phases in which a larger impact on FIP production is expected.
@@ 2.6.1 The virial theorem
Making use of Eq. (3), let us rewrite the hydrostatic equilibrium equation [Eq. (2)]
@@ 2.6.4 H and He burning
In low-mass stars ($M<2~{}{\rm M}_{\odot}$) electron degeneracy develops in the He-rich core. This occurrence has two major effects. Firstly the electron conductivity increases, contributing to an efficient energy redistribution within the core. As a result, an almost isothermal temperature profile develops within the core. On top of that, a plasma neutrino emission is more effective in the innermost layers, where the density is higher. As a consequence of this neutrino cooling, the maximum temperature moves out of the center (see Fig. 4). Then, an off-center He ignition occurs, which causes a thermonuclear runaway known as He-flash. In practice, the release of nuclear energy by the triple-$\alpha$ reactions causes an increase of the temperature not counterbalanced by an increase of the pressure. This is a consequence of the degenerate equation of state, for which the electron pressure mainly depends on their density, while it weakly depends on $T$. As shown in Fig. 4, up to $10^{10}$ $L_{\odot}$ of nuclear power are attained by the first and most powerful flash, which is followed by a series of weaker thermonuclear runaways, until the electron degeneracy is removed in the whole core. This occurrence marks the beginning of the quiescent He burning and it is called the Zero Age Horizontal Branch (ZAHB).
@@ 2.6.6 The final destiny of low-mass and intermediate-mass stars
During this phase, the star climbs the so-called AGB. The compact C-O core is surrounded by a thin He-rich mantle and a more external and very expanded H-rich envelope. The stellar radius becomes very large, i.e., a few hundreds R. Two burning shells are alternatively actives. For most of the time the shell-H burning provides enough energy to replace the energy lost by radiation. Meanwhile the He-rich mantel is accreted with the ashes of the H-burning, until it reaches a critical mass. Then, recursive thermal instability powered by the periodic activation of the He-burning takes place (thermal pulses). The discovery of these thermal instabilities [101] had a great relevance in our understanding of the chemical evolution of the Universe. About half of the elements heavier than iron are produced by low-mass AGB stars through a slow-neutron-capture nucleosynthesis process or s process [102]. Moreover, AGB stars are among the major sources of dust in a galaxy like the Milky Way. The AGB mass loss rate may be as large as $10^{-5}$ M/yr, and the wind from these stars may be as fast as $10^{6}$ cm/s. Owing to this huge mass loss, the H-rich envelope is rapidly eroded until a bare core remains. These stars end their life as C-O WD.
Figure 10: Evolution of the central temperature and density for stars having masses between 1 and 9 ${\rm M}_{\odot}$ and nearly solar composition [97]. The three principal evolutionary phases are labeled, namely: core-H burning (H), core-He burning (He) and the AGB. Note, in particular, the expansion occurring after the He ignition, as due to the release of nuclear energy, and the cooling of the core during the final part of the evolutionary tracks, when stars climb the AGB. This occurrence is a consequence of the production of plasma-neutrinos, which extracts energy from the stellar core. The red-dashed curves represent constant degeneracy parameter lines, namely: $\phi=\frac{\mu}{kT}=-10,0,10,100$, left to right. Here $\mu$ is the electron chemical potential that, in the $T=0$ limit, coincides with the Fermi energy. These curves have been calculated assuming an electron fraction $Y_{e}=0.5$, which is appropriate for the He-rich core of RGB stars as well as for the C-O core of AGB stars. Low-mass stars ($M<2.5~{}{\rm M}_{\odot}$) develop a partial degeneracy of the electron ($\phi>0$) during the first ascent of the RGB. Later on, during the AGB phase, the electron degeneracy develops again, in all stars with mass lower than $\sim 9~{}{\rm M}_{\odot}$. Eventually, the yellow area delimits the region where electrons are degenerate and relativistic. Note that the more massive AGB stars ($M>4~{}{\rm M}_{\odot}$) enter in this region when the plasma-neutrino cooling is more efficient.
@@ 3.1 Axions and axion-like particles
Several extensions of the SM of particle physics predict a hidden sector of new particles interacting very weakly with SM particles. Often, such scenarios include very light particles, for example as Nambu-Goldstone bosons of global symmetries [10, 9, 103, 104, 105, 106], or as modes associated with compactified extra-dimensions [107, 108, 109]. The study of these additional light modes goes under the name of the low-energy frontier of particle physics [18] and has become one of the most studied branches of modern particle physics. Indeed, a general interest in the FIPs has emerged in recent years, particularly since they offer explanations to several important cosmological and astrophysical problems (see Refs. [17, 16] for recent reviews).
@@ 00 KSVZ
Although different models can provide very different phenomenological results, in general, axions are highly constrained by the requirement to solve the strong CP problem (see Ref. [125] for more details). Here we just summarize a few aspects  [^7] The interested reader may find more information in Ref. [50].:
@@
Figure 12: Contours of the axion energy-loss rates per unit mass, $\varepsilon_{a}$, in erg g-1s-1, for a pure He plasma for ordinary stars, while only nucleons and electron fraction $Y_{e}=0.3$ for supernovae (SNe) and neutron stars (NSs). Different lines correspond to different channels, as shown in the legend. The Primakoff process (blue lines) is calculated for $g_{a\gamma}=10^{-11}$eV-1. The Bremsstrahlung (black lines) and Compton (green lines) processes are calculated for $g_{ae}=10^{-13}$. The onset of the degeneracy region is visible in the bending of the bremsstrahlung contours. The shaded regions are indicative of the central temperature and density of the Sun [143], Red Giant Branch (RGB) stars, horizontal branch (HB) stars and SNe are also shown. In the case of HB and RGB, these are the results of a numerical simulation of a 0.8 $M_{\odot}$ model as obtained with the FuNS code [89]. With the same code also the Betelgeuse model is obtained [26]. Regarding degenerate stars, we show the properties of WD L19-2 and two NSs, representing the case of a young NS, J172054 with an age of $\sim 650$ yr, and an old one, J1856 with an age of $4\times 10^{5}$ yr [36]. The thickness of the lines has no significance, except for SNe, where roughly represents the plausible range of density and temperature.